# The Milky Way in context: The formation of galactic discs and chemical sequences from a cosmological perspective

### v1

@@ 3.3.4 The shape of chemical sequences
Figure 13: The chemical abundance planes for a selection of auriga galaxies, showing only disc stars that were born in the Solar region ($96<R_{\rm G}/\rm{kpc}<$ 9). The histograms are coloured by the gradient of the local SFR associated with each star particle, evaluated over a time period of $\tau_{\rm birth}\pm 1$ Gyr. We calculate the SFR for stars that were born in the Solar region rather than stars found in the Solar region at $z=0$. This is so that the gradient of the SFR better reflects the birth environment of each star particle, which may change due to radial mixing and migration. The stellar mass distribution is overlaid with black contours. The formatting is otherwise as in Figure 1. Upper panels: Examples of galaxies with a flat distribution in their $\alpha$-sequences. Lower panels: Examples where there is a downward slanted distribution in their $\alpha$-sequences. Additional plots for the full suite are included online.
@@
We attribute the origin of this relation to long-lived trends in the SFR. The $\alpha$-elements are produced by Type-II SNe of short-lived massive stars, which rapidly enrich the ISM. Iron is mainly produced by the Type-Ia SNe of longer-lived binary stars, and can take $\mathcal{O}1\,$Gyr to enrich the ISM. Therefore, an increasing SFR will raise or maintain the [Mg/Fe] ratio because the iron enrichment is lagging behind the $\alpha$-enrichment. A SFR that is decreasing or unchanging over timescales of a Gyr or more will drive the same ratio down. Therefore, chemical sequences with steep downward ‘slants’ should correlate with an unchanging or decreasing SFR.
We provide evidence that this occurs in auriga in Figure 13, which shows the mean gradient of the radially-binned SFR in the chemical abundance plane for a selection of realisations. Here, we are calculating the SFR for stars that were born in the range $96<R_{\rm G}/\rm{kpc}<$ 9, though the same results are recovered when considering stars found in the same radial range at $z=0$. The upper five panels show examples where the $\alpha$-sequences are flat. There, we see that the SFR is rising for both high-$\alpha$ and low-$\alpha$ sequences. For realisations with bi-modal chemistry (see the pink coloured markers in the top-left), we can also see that the SFR declines for the under-populated region between the two sequences. This is similar to the result shown in Figure 6.
The lower five panels show examples where the $\alpha$-sequences are ‘slanted’ from top-left to bottom-right (an inverse relation between [Fe/H] and [Mg/Fe]). Each example has an initially growing SFR at high-$\alpha$, corresponding to the growth of the proto-galaxy at early times. Then, the SFR is either unchanging or decreasing throughout the formation of the low-$\alpha$ sequence. These trends are consistent across the rest of the auriga suite and for other radial ranges not shown here. This behaviour is the same as is reported in the eagle simulation suite by Mason et al. (2024), in which the formation of “knees” in the [$\alpha$/Fe] versus [Fe/H] plane are attributed to a sustained decline in the SFR.
@@ 3.3.5 Radial migration
A potential complication in the formation of galactic chemical sequences is the tendency for stars to migrate away from their initial birth sites (radial migration). This is caused by a change in their orbital angular momenta, originating from interactions with non-axisymmetric disc structure (e.g. spiral arms Sellwood & Binney, 2002; Roškar et al., 2008, central bars Minchev & Famaey, 2010; Minchev et al., 2011, or mergers and subhaloes Minchev et al., 2014b; Carr et al., 2022).
The impact of radial migration on the auriga level-3 simulations was previously discussed in Grand et al. (2018), where it was found to have little importance for the development of chemical sequences. This is consistent with idealised simulations that show radial migration is not the dominant factor with respect to the bi-modality (Minchev et al., 2014a). For completeness, we have performed our own analysis on all of the level-4 simulations, for which we show just one example in Figure 14.
This figure is generally representative of the rest of the suite, and exhibits a few key features. Firstly, there is very little bulk migration occurring for the high-$\alpha$ and low-$\alpha$ sequences around the Solar neighbourhood ($96<R_{\rm G}/\rm{kpc}<$ 9). This shows that the bi-modality at these radii does not arise through the radial migration of stars, but specifically from the gas chemistry and SFR local to the Solar neighbourhood.
@@
This result is qualitatively in agreement with the broadening of [Fe/H] distributions reported in idealised simulations of a MW-mass galaxy in Grand et al. (2015), which was also caused by radial migration processes. Similarly, Haywood et al. (2013) use observational data to argue that radial migration would not have significantly contaminated the Solar neighbourhood in the MW.
@@ 4.1 A consecutive formation of the chemical sequences
We have found that, in all cases except for one unusual outlier (Au-1), the formation of $\alpha$-sequences is sequential. Any apparent temporal overlap between two sequences is due to the spread in chemical abundances for mono-age populations rather than a faithful coeval formation. This, then, is accurate to first-order to what is observed in the MW, with many studies finding that the high-$\alpha$ sequence formed rapidly before $\sim 8\,$Gyr ago and the low-$\alpha$ sequence formed over a much longer period of time thereafter (e.g. Fuhrmann, 1998; Bensby et al., 2005; Fuhrmann, 2011; Haywood et al., 2013; Mackereth et al., 2017; Hayden et al., 2017).
@@
However, there is also observational evidence for a fraction of low-$\alpha$ stars that formed at lookback times older than 10 Gyr ago, in consort with the high-$\alpha$ sequence (e.g. Beraldo e Silva et al., 2021; Gent et al., 2024; Borbolato et al., 2025). The authors in Gent et al. (2024) show that up to 20 per cent of stars in the low-$\alpha$ sequence formed prior to 9 Gyr ago using Gaia-ESO data for main sequence turnoff stars (MSTOs) and subgiants. Similarly, Borbolato et al. (2025) use the StarHorse code (Santiago et al., 2016) to determine stellar ages for Solar neighbourhood stars in both APOGEE and LAMOST spectroscopic survey data, finding the fraction of low-$\alpha$ stars with ages older than 11 Gyrs to be over a third in the case of APOGEE and nearly a half in the case of LAMOST. These suggest the early Galaxy was very capable of achieving low-$\alpha$ chemistry even before the accretion of the GSE at around 10 Gyr ago.
@@
One possible explanation is that these stars were born within intermediate-mass gas clumps that are predisposed to form at high redshifts when gas fractions are higher (see Bournaud et al., 2009; Clarke et al., 2019). The sub-grid physics model in auriga does not resolve the formation of such gas clumps, and further studies would be needed to explore the impact of these physics in fully cosmological simulations.
@@
Observational studies find that the high-$\alpha$ disc transitioned to almost exclusively low-$\alpha$ formation very rapidly at approximately 8-9 Gyr ago (Linden et al., 2017; Sahlholdt et al., 2022; Borbolato et al., 2025). This, then, is just 1-2 Gyr after the likely accretion time of the GSE merger (e.g. Belokurov et al., 2018; Helmi et al., 2018; Mackereth et al., 2019; Kruijssen et al., 2019). However, these investigations are often restricted to the Solar neighbourhood and/or limited stellar populations (e.g. see the review of Soderblom, 2010), and by high uncertainty in stellar age determinations. Likewise, the exact accretion time of the GSE is still uncertain to within a precision of 1-2 Gyrs. In this work, we find that gas from GSE-like mergers was not instrumental for the formation low-$\alpha$ sequences via their gas donation or metallicity dilution, but could still have been important in other respects. For example, a massive injection of gas from a singular direction establishes a high angular momentum in the gas, and may also help to align the disc with the angular momentum of wider-scale CGM gas, which ultimately aids in the build-up of an extended stellar disc at higher radii (e.g. see the investigation into Au-18 in Merrow et al., 2024).
@@
To better understand the timeline of disc formation, it will be crucial to obtain stellar age estimates with higher precisions and for wider samples. This will be addressed by upcoming asteroseismology on thousands of stars in ESA’s half-sky PLATO survey (Rauer et al., 2025), or for a million stars in the further future by the all-sky Chronos mission (Michel et al., 2021).
@@ 4.2.1 Gas fuelling and dilution
The GSE would be an opportune vehicle for the donation of fresh metal-poor gas into the ancient Galaxy, as called for by the “two-infall” model of Chiappini et al. (1997, 2001). This could reset the gas metallicity to low values whilst also providing the fuel to support the formation of the low-$\alpha$ sequence. The chemical evolution models in Spitoni et al. (2019, 2021, 2023) show that a metallicity dilution of approximately a full dex is able to reproduce the low-$\alpha$ sequence in the MW.
@@
In our investigation, we have found that singular merger events do not induce a long-lived and significant ($\Delta[\rm{Fe/H}]>0.5$) dilution in the gas discs of their hosts. The metal-poor gases donated by minor mergers are rapidly mixed into the surrounding CGM with low impact on the overall metallicity, whereas major mergers increasingly approach parity with the metallicity of their host and so rarely yield a dilution greater than $\sim 0.3\,$dex. In fact, we find some realisations with GSE-like mergers that exhibit no metallicity overlap between their two $\alpha$-sequences, or do not have multiple sequences at all. Furthermore, the gas from most mergers forms only a minor component of subsequent in-situ disc stars (with massive major mergers of mass ratio $>1:2$ being an exception).
@@
This is not to suggest that mergers are entirely incapable of inducing dilution. For instance, Grand et al. (2018) demonstrated that dilution occurs at the edge of the gas disc around the time of a late gas-rich merger in the high-resolution version of Au-23. In the version of Au-23 used in this study, we observe that dilution begins prior to the arrival of the merger gas, suggesting that the merger’s gravitational influence may draw metal-poor gas from the CGM onto the disc’s outskirts (which is then supplemented by the merger gas itself), thereby indirectly causing dilution. Nonetheless, this dilution is insufficient to account for the metallicity overlap observed in the Milky Way’s chemical sequences, and the properties of this merger event are inconsistent with those expected for the GSE.
@@
We show that the greatest metal dilution occurs when metal-poor merger gases mix with the highly-enriched central gas of the host, in agreement with Buck (2020) who find that merger-induced metal dilution in the nihao simulations was maximised in the inner 3 kpc. Simulations have also shown that gravitational torques from satellites can stimulate the inflow of metal-poor CGM gases into galactic centres (Di Matteo et al., 2007; Bustamante et al., 2018; Moreno et al., 2021; Faria et al., 2025), contributing to a central starburst and further metallicity dilution. Yet, observations of the MW indicate that the metallicity overlap between the two $\alpha$-sequences is least pronounced from within $\sim 6\,$kpc (e.g. Queiroz et al., 2020). Therefore, whilst these effects may have occurred in the MW, they were clearly not the primary drivers in the development of the chemical sequences. Then, it is unlikely that the metallicity overlap in the chemical sequences of the MW was caused by a highly radial infall such as the GSE, if it was caused by a merger event at all.
Contrary to our interpretations, Buck (2020) argue that metallicity dilution from gas-rich minor mergers (of less than $1:10$ mass ratio) explains the overlapping metallicities of bi-modal chemical sequences. However, we suggest that the metal dilution following those mergers is insufficient. The metallicity overlap is as great as 0.5 dex (see their figure 1). Yet, from inspection of their figure 7, the metal dilution following gas-rich mergers is typically far less than half of this, and is short-lived in most cases. Then, this would be consistent with the findings presented in this work. In fact, our result is general and should hold for any reasonable galaxy formation model in a $\Lambda$CDM Universe. In Appendix A we demonstrate this further using simple arguments involving empirical relations (e.g. stellar halo mass relation Moster et al., 2013) and observations (e.g. the stellar mass-metallicity relation Kirby et al., 2013).
@@ 4.2.2 Influence on the star formation rate
In Section 3.2, we showed that auriga galaxies with particularly pronounced bi-modal chemistry exhibit a lull in their star formation following merger-driven, centrally concentrated starbursts. Indeed, this is a behaviour also seen among other cosmological simulations like simba (Rodríguez Montero et al., 2019), IllustrisTNG (Hani et al., 2020), vintergatan (Segovia Otero et al., 2022), and fire-2 (Li et al., 2025). In auriga, the enhanced SFR has either fully or partially depleted the star-forming gas, ultimately leading to a temporary suppression of the star formation. This is reflected in observational data from SDSS finding evidence that starbursts correlate with features of merger interactions (Luo et al., 2014). This is a possible scenario for the MW, where the two $\alpha$-sequences are divided by an under-populated gap, and there is a clear star formation quiescence at $10\text{-}8\,$Gyr ago (e.g. Haywood et al. 2016, 2018; Spitoni et al. 2024, and see also figure 7 in Snaith et al. 2015).
Our finding is consistent with Beane (2025), where idealised simulations show that a brief star formation hiatus of a few hundred Myr could arise due to gas starvation following a GSE-like merger (though with great sensitivity to the infall inclination of said merger). The hiatus does not necessarily need to owe to the influence of a merger, and in a follow-up work they propose AGN-induced quenching coincident with the formation of a stellar bar as another possible mechanism (Beane et al., 2024, and see also Haywood et al., 2018). Additionally, Grand et al. (2018) demonstrate than the time-evolution in the size of the gas disc can lead to the formation of a chemical bi-modality over the pertinent radii. Disc-galaxy evolution models in Noguchi (2018) also predict that a star formation hiatus of approximately 2 Gyr can create a bi-modality. Finally, simulations in Lian et al. (2020) predict a starburst followed by rapid quenching could be responsible for forming the thick disc. These are compatible with observations of the MW from APOGEE, which finds evidence of a $\sim 2\,$Gyr quenching period roughly 9 Gyr ago (see Haywood et al., 2016).

### v2

@@ 3.3.4 The shape of chemical sequences
Figure 13: The chemical abundance planes for a selection of auriga galaxies, showing only disc stars that were born in the Solar region ($6<R_{\rm G}/\rm{kpc}<9$). The histograms are coloured by the gradient of the local SFR associated with each star particle, evaluated over a time period of $\tau_{\rm birth}\pm 1$ Gyr. We calculate the SFR for stars that were born in the Solar region rather than stars found in the Solar region at $z=0$. This is so that the gradient of the SFR better reflects the birth environment of each star particle, which may change due to radial mixing and migration. The stellar mass distribution is overlaid with black contours. The formatting is otherwise as in Figure 1. Upper panels: Examples of galaxies with a flat distribution in their $\alpha$-sequences. Lower panels: Examples where there is a downward slanted distribution in their $\alpha$-sequences. Additional plots for the full suite are included online.
@@
We attribute the origin of this relation to long-lived trends in the SFR. The $\alpha$-elements are produced by Type-II SNe of short-lived massive stars, which rapidly enrich the ISM. Iron is mainly produced by the Type-Ia SNe of longer-lived binary stars, and can take $\sim 1\,$Gyr to enrich the ISM. Therefore, an increasing SFR will raise or maintain the [Mg/Fe] ratio because the iron enrichment is lagging behind the $\alpha$-enrichment. A SFR that is decreasing or unchanging over timescales of a Gyr or more will drive the same ratio down. Therefore, chemical sequences with steep downward ‘slants’ should correlate with an unchanging or decreasing SFR.
We provide evidence that this occurs in auriga in Figure 13, which shows the mean gradient of the radially-binned SFR in the chemical abundance plane for a selection of realisations. Here, we are calculating the SFR for stars that were born in the range $6<R_{\rm G}/\rm{kpc}<9$, though the same results are recovered when considering stars found in the same radial range at $z=0$. The upper five panels show examples where the $\alpha$-sequences are flat. There, we see that the SFR is rising for both high-$\alpha$ and low-$\alpha$ sequences. For realisations with bi-modal chemistry (see the pink coloured markers in the top-left), we can also see that the SFR declines for the under-populated region between the two sequences. This is similar to the result shown in Figure 6.
The lower five panels show examples where the $\alpha$-sequences are ‘slanted’ from top-left to bottom-right (an inverse relation between [Fe/H] and [Mg/Fe]). Each example has an initially growing SFR at high-$\alpha$, corresponding to the growth of the proto-galaxy at early times. Then, the SFR is either unchanging or decreasing throughout the formation of the low-$\alpha$ sequence. These trends are consistent across the rest of the auriga suite and for other radial ranges not shown here. This behaviour is the same as is reported in the eagle simulation suite by mason2024, in which the formation of “knees” in the $\alpha$/Fe] versus [Fe/H] plane are attributed to a sustained decline in the SFR.
@@ 3.3.5 Radial migration
A potential complication in the formation of galactic chemical sequences is the tendency for stars to migrate away from their initial birth sites (radial migration). This is caused by a change in their orbital angular momenta, originating from interactions with non-axisymmetric disc structure (e.g. spiral arms sellwood2002; rovskar2008, central bars minchev2010; minchev2011, or mergers and subhaloes minchev2014b; carr2022).
The impact of radial migration on the auriga level-3 simulations was previously discussed in grand2018, where it was found to have little importance for the development of chemical sequences. This is consistent with idealised simulations that show radial migration is not the dominant factor with respect to the bi-modality (minchev2014). For completeness, we have performed our own analysis on all of the level-4 simulations, for which we show just one example in Figure 14.
This figure is generally representative of the rest of the suite, and exhibits a few key features. Firstly, there is very little bulk migration occurring for the high-$\alpha$ and low-$\alpha$ sequences around the Solar neighbourhood ($6<R_{\rm G}/\rm{kpc}<9$). This shows that the bi-modality at these radii does not arise through the radial migration of stars, but specifically from the gas chemistry and SFR local to the Solar neighbourhood.
@@
This result is qualitatively in agreement with the broadening of [Fe/H] distributions reported in idealised simulations of a MW-mass galaxy in grand2015, which was also caused by radial migration processes. Similarly, haywood2013 use observational data to argue that radial migration would not have significantly contaminated the Solar neighbourhood in the MW.
@@ 4.1 A consecutive formation of the chemical sequences
We have found that, in all cases except for one unusual outlier (Au-1), the formation of $\alpha$-sequences is sequential. Any apparent temporal overlap between two sequences is due to the spread in chemical abundances for mono-age populations rather than a faithful coeval formation. This, then, is accurate to first-order to what is observed in the MW, with many studies finding that the high-$\alpha$ sequence formed rapidly before $\sim 8\,$Gyr ago and the low-$\alpha$ sequence formed over a much longer period of time thereafter (e.g. fuhrmann1998; bensby2005; fuhrmann2011; haywood2013; mackereth2017; hayden2017).
@@
However, there is also observational evidence for a fraction of low-$\alpha$ stars that formed at lookback times older than 10 Gyr ago, in consort with the high-$\alpha$ sequence (e.g. beraldo2021; gent2024; borbolato2025). The authors in gent2024 show that up to 20 per cent of stars in the low-$\alpha$ sequence formed prior to 9 Gyr ago using Gaia-ESO data for main sequence turnoff stars (MSTOs) and subgiants. Similarly, borbolato2025 use the StarHorse code (santiago2016) to determine stellar ages for Solar neighbourhood stars in both APOGEE and LAMOST spectroscopic survey data, finding the fraction of low-$\alpha$ stars with ages older than 11 Gyrs to be over a third in the case of APOGEE and nearly a half in the case of LAMOST. These suggest the early Galaxy was very capable of achieving low-$\alpha$ chemistry even before the accretion of the GSE at around 10 Gyr ago.
@@
One possible explanation is that these stars were born within intermediate-mass gas clumps that are predisposed to form at high redshifts when gas fractions are higher (see bournaud2009; clarke2019). The sub-grid physics model in auriga does not resolve the formation of such gas clumps, and further studies would be needed to explore the impact of these physics in fully cosmological simulations.
@@
Observational studies find that the high-$\alpha$ disc transitioned to almost exclusively low-$\alpha$ formation very rapidly at approximately 8-9 Gyr ago (linden2017; sahlholdt2022; borbolato2025). This, then, is just 1-2 Gyr after the likely accretion time of the GSE merger (e.g. belokurov2018; helmi2018; mackereth2019; Kruijssen2019). However, these investigations are often restricted to the Solar neighbourhood and/or limited stellar populations (e.g. see the review of soderblom2010), and by high uncertainty in stellar age determinations. Likewise, the exact accretion time of the GSE is still uncertain to within a precision of 1-2 Gyrs. In this work, we find that gas from GSE-like mergers was not instrumental for the formation low-$\alpha$ sequences via their gas donation or metallicity dilution, but could still have been important in other respects. For example, a massive injection of gas from a singular direction establishes a high angular momentum in the gas, and may also help to align the disc with the angular momentum of wider-scale CGM gas, which ultimately aids in the build-up of an extended stellar disc at higher radii (e.g. see the investigation into Au-18 in merrow2024).
@@
To better understand the timeline of disc formation, it will be crucial to obtain stellar age estimates with higher precisions and for wider samples. This will be addressed by upcoming asteroseismology on thousands of stars in ESA’s half-sky PLATO survey (PLATO), or for a million stars in the further future by the all-sky Chronos mission (chronos).
@@ 4.2.1 Gas fuelling and dilution
The GSE would be an opportune vehicle for the donation of fresh metal-poor gas into the ancient Galaxy, as called for by the “two-infall” model of chiappini1997; chiappini2001. This could reset the gas metallicity to low values whilst also providing the fuel to support the formation of the low-$\alpha$ sequence. The chemical evolution models in spitoni2019; spitoni2021; spitoni2023 show that a metallicity dilution of approximately a full dex is able to reproduce the low-$\alpha$ sequence in the MW. Improving stellar age estimates and extensions to existing catalogues have allowed for more detailed chemical evolution models, highlighting some features of the low-$\alpha$ disc that are difficult to explain with a two-infall model (see dubay2025).
@@
In our investigation, we have found that singular merger events do not induce a long-lived and significant ($\Delta[\rm{Fe/H}]>0.5$) dilution in the gas discs of their hosts. The metal-poor gases donated by minor mergers are rapidly mixed into the surrounding CGM with low impact on the overall metallicity, whereas major mergers increasingly approach parity with the metallicity of their host and so rarely yield a dilution greater than $\sim 0.3\,$dex. In fact, we find some realisations with GSE-like mergers that exhibit no metallicity overlap between their two $\alpha$-sequences, or do not have multiple sequences at all. Furthermore, the gas from most mergers forms only a minor component of subsequent in-situ disc stars (with massive major mergers of mass ratio >1:2>1:2 being an exception).
@@
This is not to suggest that mergers are entirely incapable of inducing dilution. For instance, grand2018 demonstrated that dilution occurs at the edge of the gas disc around the time of a late gas-rich merger in the high-resolution version of Au-23. In the version of Au-23 used in this study, we observe that dilution begins prior to the arrival of the merger gas, suggesting that the merger’s gravitational influence may draw metal-poor gas from the CGM onto the disc’s outskirts (which is then supplemented by the merger gas itself), thereby indirectly causing dilution. Nonetheless, this dilution is insufficient to account for the metallicity overlap observed in the Milky Way’s chemical sequences, and the properties of this merger event are inconsistent with those expected for the GSE.
@@
We show that the greatest metal dilution occurs when metal-poor merger gases mix with the highly-enriched central gas of the host, in agreement with buck2020 who find that merger-induced metal dilution in the nihao simulations was maximised in the inner 3 kpc. Simulations have also shown that gravitational torques from satellites can stimulate the inflow of metal-poor CGM gases into galactic centres (dimatteo2007; bustamante2018; moreno2021; faria2025), contributing to a central starburst and further metallicity dilution. Yet, observations of the MW indicate that the metallicity overlap between the two $\alpha$-sequences is least pronounced from within $\sim 6\,$kpc (e.g. queiroz2020). Therefore, whilst these effects may have occurred in the MW, they were clearly not the primary drivers in the development of the chemical sequences. Then, it is unlikely that the metallicity overlap in the chemical sequences of the MW was caused by a highly radial infall such as the GSE, if it was caused by a merger event at all.
[…]
@@ 4.2.2 Influence on the star formation rate
In Section 3.2, we showed that auriga galaxies with particularly pronounced bi-modal chemistry exhibit a lull in their star formation following merger-driven, centrally concentrated starbursts. Indeed, this is a behaviour also seen among other cosmological simulations like simba (rodriguez2019), IllustrisTNG (hani2020), vintergatan (segovia2022), and fire-2 (li2025). In auriga, the enhanced SFR has either fully or partially depleted the star-forming gas, ultimately leading to a temporary suppression of the star formation. This is reflected in observational data from SDSS (the Sloan Digital Sky Survey) finding evidence that starbursts correlate with features of merger interactions (luo2014). This is a possible scenario for the MW, where the two $\alpha$-sequences are divided by an under-populated gap, and there is a clear star formation quiescence at $10\text{-}8\,$Gyr ago (e.g. haywood2016; haywood2018; spitoni2024, and see also figure 7 in snaith2015).
Our finding is consistent with beane2024a, where idealised simulations show that a brief star formation hiatus of a few hundred Myr could arise due to gas starvation following a GSE-like merger (though with great sensitivity to the infall inclination of said merger). The hiatus does not necessarily need to owe to the influence of a merger, and in a follow-up work they propose AGN-induced quenching coincident with the formation of a stellar bar as another possible mechanism (beane2024b, and see also haywood2018). Additionally, grand2018 demonstrate than the time-evolution in the size of the gas disc can lead to the formation of a chemical bi-modality over the pertinent radii. Disc-galaxy evolution models in noguchi2018 also predict that a star formation hiatus of approximately 2 Gyr can create a bi-modality. Finally, simulations in lian2020 predict a starburst followed by rapid quenching could be responsible for forming the thick disc. These are compatible with observations of the MW from APOGEE, which finds evidence of a $\sim 2\,$Gyr quenching period roughly 9 Gyr ago (see haywood2016).